OpenStudy (samara8954):

Alan can run 6 miles in 48 minutes. If he maintains the same pace, what is the greatest number of miles he can run in 192 minutes? 24 48 72 36

OpenStudy (samara8954):

@Da_Homie

OpenStudy (anonymous):

A

jimthompson5910 (jim_thompson5910):

(6 miles)/(48 minutes) = (x miles)/(192 minutes) 6/48 = x/192 6*192 = 48*x ... cross multiply x = ??

OpenStudy (anonymous):

i would say a and i like your response jim_thompson5910